A parabolic satellite television antenna has a diameter of 6 feet and is 1.6 feet deep. How far is the focus from the vertex?
I have tried every formula that I can think of to get this one to work out correctly for me and I'm coming up short.

Please help.
Thank you in advance

Answer by stanbon(48580)

A parabolic satellite television antenna has a diameter of 6 feet and is 1.6 feet deep. How far is the focus from the vertex?
----------------------------
Draw the parabola on an xy coordinate grill.
Put the vertex at (0,1.6)
Then the parabola will meet the x-axis at (-3,0) and (3,0)
----------------
So the equation of the parabola will be y = -x^2 + 1.6
or x^2 = -y - 1.6
or x^2 = -(y+1.6)
-----------------
This has the form x^2 = 4p(y-k), so 4p = -1 and p = -1/4
---
The distance from the vertex to the focus is p so the
focus is at y = 1.6-(1/4) = 1.6-0.25 = 1.375
-------------------
Ans: the focus is (1/4) ft. from the vertex.
======================
Cheers,
Stan H.